### v1

# GraphThinker: Reinforcing Video Reasoning with Event Graph Thinking

Abstract
Video reasoning requires understanding the causal relationships between events in a video. However, such relationships are often implicit and costly to annotate manually. While existing multimodal large language models (MLLMs) often infer event relations through dense captions or video summaries for video reasoning, such modeling still lacks causal understanding. Without explicit causal structure modeling within and across video events, these models suffer from hallucinations during the video reasoning. In this work, we propose GraphThinker, a reinforcement finetuning-based method that constructs structural event-level scene graphs and enhances visual grounding to jointly reduce hallucinations in video reasoning. Specifically, we first employ an MLLM to construct an event-based video scene graph (EVSG) that explicitly models both intra- and inter-event relations, and incorporate these formed scene graphs into the MLLM as an intermediate thinking process. We also introduce a visual attention reward during reinforcement finetuning, which strengthens video grounding and further mitigates hallucinations. We evaluate GraphThinker on two datasets, RexTime and VidHalluc, where it shows superior ability to capture object and event relations with more precise event localization, reducing hallucinations in video reasoning compared to prior methods.
1 Introduction
Video reasoning requires understanding the causal relationships between events in a video to answer complex natural-language queries [4, 8, 11, 21, 34, 73, 14]. Unlike image-level understanding, it demands reasoning over temporal dependencies, event transitions, and cross-event causality. Such capabilities are crucial for high-level applications, including instructional video understanding [2, 61, 69], embodied decision-making [67, 36, 15], and assistive AI systems [17, 60], that rely on reliable video reasoning.
Figure 1: Current MLLMs [2] implicitly model event relations in video by token correlations. This often leads to hallucinations in video reasoning. For instance, when determining the temporal order of events, QwenVL-2.5 tends to yield temporal hallucinations due to a lack of explicit event-level evidence validation and reasoning. In contrast, our GraphThinker explicitly models both intra- and inter-event relations in accordance with selective visual attentions in a video to constrain reasoning under these structured event relations anchored by fine-grained visual evidence, reducing hallucinations and improving temporal consistency in reasoning.
Recent advances in multimodal large language models (MLLMs) have substantially improved video reasoning by integrating visual and linguistic understanding. Such models commonly are enhanced by post-training based supervised fine-tuning (SFT) or reinforcement fine-tuning (RFT) for their video reasoning. Among them, SFT-based approaches  [26, 19, 72, 71, 21, 16, 55] teach models to follow annotated reasoning paths, while RL-based methods [13, 62, 33, 57, 6, 38, 23, 44] optimize the models to maximize formalized reward signals. Despite effectiveness, the resulting models are often fluent at reasoning generation but weak at grounding true visual cues in the input video, inducing hallucinations in their reasoning [43, 45, 29, 10, 3].
To incorporate more event relations within a video into the reasoning process, some works [64, 35, 31, 16, 21, 54] employ text-based assistance, such as dense captions or key-frame summaries, that capture rough sequential information in a video. However, such captions or summaries are overly redundant and inherently unstructured, making the model prone to hallucinations and hindering its ability to extract useful information for reasoning. Others [12, 27, 49] use video-level scene graphs that mainly capture global object relations, but these video-level graphs fail to represent the dynamic dependencies and interactions among events within a video. As a result, current video scene graphs lack an explicit event-level relational structure, causing MLLMs confusion in event ordering and causal reasoning and leading to hallucinations in their reasoning (as shown in Fig. 1).
To minimize hallucinations during video reasoning, in this work, we construct fine-grained, temporally grounded event-level scene graphs to guide video reasoning of an MLLM. To this end, we introduce GraphThinker, a reinforcement finetuning-based method that constructs structural event-level scene graphs and enhances visual grounding to jointly reduce hallucinations in video reasoning.
Specifically, we construct structured event-based video scene graphs (EVSGs) that explicitly encode fine-grained event relations to guide the MLLM’s reasoning. Our EVSGs are constructed without human annotations by a self-generate and self-refine process. We first use an MLLM to generate multi-grained event-level dense captions, and then prompt the MLLM to refine the generated captions into structured scene graphs. Each graph captures intra-event relations through object nodes and their semantic interactions within an event, while inter-event relations are modeled as timestamp-based temporal edges connecting event subgraphs. Such graphs provide temporal event relations that constrain reasoning and form an intermediate thinking process, thus mitigating hallucinations.
Moreover, we further enhance the video grounding capability of the exploited MLLM through reinforcement finetuning. We introduce a visual attention reward that encourages the MLLM to increase its visual attention scores during video reasoning. This visual attention promotion guides the MLLM to actively explore visual cues for reasoning rather than mainly relying on the input EVSGs. Together, the formed textual EVSGs and visual attention reward provide complementary benefits that enable MLLMs to achieve visually grounded, temporally consistent, and hallucination-mitigated video reasoning.
We summarize our contributions as follows:
  • We observe that existing video MLLMs lack the explicit modeling of event relations within input videos, which limits their ability to perform coherent and causally consistent reasoning, thus introducing hallucinations.
  • We propose GraphThinker, a reinforcement fine-tuning method for modeling explicit causal and event-level reasoning in videos. GraphThinker jointly integrates a structured textual graph and a visual reward to enforce visual grounding, mitigating hallucinations in video reasoning.
  • GraphThinker achieves effective improvements over existing state-of-the-art video reasoning MLLMs on two video reasoning benchmarks, RexTime and VidHalluc.
2 Related Work
2.1 Video Scene Graph Generation
Video scene graphs (VSGs) decompose videos into objects and their pairwise relations to model how objects interact over time [28, 68, 63]. These structured representations bridge low-level perception and high-level reasoning, enabling interpretable video understanding. Traditional VSG methods include two-stage approaches [9, 50, 48], which first detect objects and then classify relations, and one-stage approaches [39], which predict entities and relations simultaneously. However, both are limited by closed vocabularies and the quality of external detectors.
Open-vocabulary VSG methods [22, 37] often use vision-language models to predict relational triplets in unconstrained semantic spaces, thereby enhancing generalisation to unseen concepts; however, they often produce visually ungrounded or temporally inconsistent graphs. Moreover, existing methods [12, 49] typically generate coarse video-level representations that overlook fine-grained temporal boundaries and event dependencies, leading to imprecise temporal understanding and potential hallucinations in video reasoning. To address these, we propose the event-based video scene graph (EVSG), which explicitly encodes object and event relations over time, providing a structured representation that enables MLLMs to perform visually grounded and temporally consistent video reasoning.
Figure 2: An overview of the GraphThinker video reasoning model. GraphThinker first employs an MLLM to generate multi-grained dense captions in sequence for a video. It deploys the same MLLM again to select keywords as graph nodes before iteratively optimizing them in constructing an event-based graph (EVSG). This EVSG is then used to serve as a fine-grained representation of structured event relations for reasoning. Given EVSG, we further develop an Event Graph-based RL post-training method with a visual attention reward in a reinforcement learning process conditioned by EVSG for selecting more attentive visual evidence. Together, GraphThinker achieves visually more grounded and temporally more consistent video reasoning.
2.2 Post-training MLLMs for Video Reasoning
Video reasoning involves identifying relevant visual evidence and inferring event-level temporal and causal relationships to answer natural language queries. Recent MLLMs address this by post-training, typically through supervised fine-tuning (SFT) or reinforcement fine-tuning (RFT). SFT-based methods employ chain-of-thought (CoT) datasets to learn reasoning over video inputs. Specifically, they [26, 19, 72, 71, 21, 16, 55] train models on annotated chain-of-thought paths, promoting structured and interpretable reasoning, but suffer from generalisation limitations in novel scenarios.
RFT-based methods [13, 62, 33, 57, 6, 38, 23, 44] instead optimise reasoning trajectories under reward signals. These methods offer greater flexibility yet often over-rely on language priors, resulting in visually ungrounded reasoning [43, 45, 29, 10, 3]. Other methods improve reasoning by visual summaries [35, 16, 21], captions [64, 31, 54], or video-level scene graphs [12, 49, 27] with additional thinking rewards [45, 10, 29], but they struggle to capture fine-grained event relations, leading to hallucinated reasoning. To address these issues, we develop an event graph-based RFT approach, with a visual attention reward that guides an MLLM to extract the most informative cues in the input video for visually grounded reasoning.
3 Method
3.1 Task Definition
We aim to train a model, typically an MLLM nowadays, for hallucination-free video reasoning, i.e., visually grounded and temporally consistent reasoning. We apply our GraphThinker on the task of video reasoning [4], also referred to as temporal video grounding [26] or grounded VQA [65]. The objective of video reasoning is to enable a model to not only reason the correct answer to a complex video question but also to accurately locate the temporal segment within the video that supports the reasoning process. Formally, given an input video $V\in\mathbb{R}^{N\times H\times W\times 3}$ and a textual question $q$, where $N$ denotes the number of frames, and $H$ and $W$ are frame dimensions, the model $\Phi$ aims to jointly predict a reasoning answer $A$ and the corresponding grounded temporal window $T=(t^{s},t^{e})$, where $t^{s}$ and $t^{e}$ denote the start and end time of the temporal window, respectively. The task can be defined as:
$T,A]=\Phi(V,q).$ (1)
Figure 3: An Example of the proposed Event-based Video Scene Graph (EVSG) for video reasoning. The EVSG is composed of event subgraphs derived from event-level captions, with each subgraph corresponding to start–end timestamps and a set of triplets representing object interactions and spatial relationships to capture intra-event semantics. event subgraphs are sequentially linked by timestamp-based edges, forming a hierarchical structure that explicitly models both intra-event and inter-event relations for temporally consistent reasoning.
3.2 Recap of GRPO
Group Relative Policy Optimization (GRPO) [18, 62, 24] is a reinforcement learning method derived from Proximal Policy Optimization (PPO) that directly optimizes the policy of a language model without requiring a critic. The goal is to learn a policy $\pi_{\theta}$ that generates accurate reasoning outputs $o$. Formally, given a video $V$ and a question $q$, the MLLM samples a group of $G$ candidate reasoning outputs $o=\{o_{1},\dots,o_{G}\}$, where $o_{i}=[T_{i},A_{i}],i\in\{1,\dots,G\}$ in video reasoning grounding task. A reward function evaluates each candidate response by comparing it against the ground-truth, producing a set of scores $r=\{r_{1},\dots,r_{G}\}$. GRPO then normalizes the rewards across the group and defines the group-level objective to maximize a weighted sum reward $R(o)$, defined as:
$R(o)=\sum_{k=1}^{G}\frac{\pi_{\theta}(o_{k})}{\pi_{\theta}^{\text{old}}(o_{k})}\cdot\frac{r_{k}-\mathrm{mean}(r)}{\mathrm{std}(r)},$ (2)
where $\pi_{\theta}^{\text{old}}$ denotes the policy before the current update. To ensure stable training and prevent excessive deviation from the original policy, a KL-divergence regularization is applied:
$\max_{\pi_{\theta}}\;\mathbb{E}_{o\sim\pi_{\theta}^{\text{old}}}\Big[R(o)-\beta\,\mathrm{KL}\big(\pi_{\theta}\;\|\;\pi_{\text{ref}}\big)\Big],$ (3)
where $\pi_{\theta}^{\text{old}}$ denotes the previous policy used for sampling candidate outputs, $\pi_{\text{ref}}$ is a reference policy to constrain excessive deviation, and $\beta$ controls the strength of the KL penalty.
However, without explicit event relation constraints in GRPO post-training, MLLMs still generate linguistically coherent yet visually inconsistent reasoning at test time. Thus, explicitly modeling event-level relationships in GRPO post-training is crucial for reducing hallucinations in video reasoning. In this work, we leverage explicit event relations representation (e.g., EVSG) to enhance the reasoning capability of Multimodal Large Language Models (MLLMs) with GRPO.
3.3 Event-based Video Scene Graph Generation
To provide explicit reasoning structure and temporal dependency cues for event causal thinking, we construct an Event-based Video Scene Graph (EVSG) for each video. The EVSG serves as a structured event relation representation that encodes spatial relations and temporal dependencies among objects and events, providing an intermediate thinking process to constrain MLLM reasoning, thus mitigate hallucinations.
Given an input video, we first employ an MLLM to generate multi-grained dense captioning for each video. To achieve it, we set different upper limits on the number of events and allow the MLLM to automatically produce captions under these constraints. Specifically, we define three event limits (e.g., 5, 10, and 15, based on the results in Tab. 4 ) and let the model segment the video accordingly into 5, 10, or 15 consecutive and non-overlapping temporal segments. For each segment, the model generates a corresponding caption. Finally, the MLLM generates dense captions at three levels of granularity: coarse, middle, and fine, denoted as $C^{\text{coarse}}$, $C^{\text{middle}}$, and $C^{\text{fine}}$, respectively. The goal of producing multi-grained dense captions is to mitigate potential hallucinations during caption generation. Cross-level comparison of coarse and fine-grained captions detects inconsistencies, suppressing hallucinated details before graph construction.
Next, we employ the MLLM to construct the Event-based Video Scene Graph using the self-generated multi-grained dense captions. The process of building the graph from captions consists of two stages: generation and refinement. In the generation stage, the model primarily uses the middle-grained captions to build the initial event graph. Guided by predefined instructions and templates, the MLLM extracts the start and end timestamps for each event from the captions. It then identifies key object interactions described in each event and converts them into a set of video scene graph triplets, following the conventional ⟨subject–relation–object⟩ format used in traditional video scene graph representations. Formally, given the middle-grained captions
$C^{middle}=\{c_{1},c_{2},...,c_{n}\},$ (4)
where each $c_{i}$ represents the textual description of the $i$-th segmented event, and $n$ denotes the number of segmented events determined by the predefined event number limit, the MLLM maps each caption $c_{i}$ to a structured event subgraph $e_{i}$ as:
$e_{i}=f_{\text{MLLM}}(c_{i})=(t_{i}^{s},t_{i}^{e},\mathcal{T}_{i}),$ (5)
, and $\mathcal{T}_{i}=\{(s_{i},r_{i},o_{i})\}$ represents the set of scene graph triplets extracted from the event description, where $(s,r,o)$ denotes a triplet of⟨subject–relation–object⟩extracted by semantic parsing and temporal alignment. The initial event-level graph $\mathcal{G}_{\text{init}}$ is then defined as:
$\mathcal{G}_{\text{init}}=\bigl(\mathcal{V},\mathcal{E}\bigr),\quad\begin{aligned} \mathcal{V}&=\{e_{1},\dots,e_{n}\},\\ \mathcal{E}&=\{\epsilon_{i,j}\mid t_{i}^{e}\leq t_{j}^{s}\}.\end{aligned}$ (6)
where $\epsilon_{i,j}\in\{0,1\}$ denotes an edge from event subgraph $e_{i}$ to $e_{j}$, present if $t_{i}^{e}\leq t_{j}^{s}$, explicitly encoding the global temporal dependencies across events.
In the second stage, we refine the initial event graph $\mathcal{G}{\text{init}}$ using both coarse- and fine-grained captions as complementary evidence. The coarse-grained captions $C^{coarse}$ offer a global narrative, while the fine-grained captions $C^{fine}$ provide detailed event descriptions. These are jointly fed with $\mathcal{G}{\text{init}}$ into the MLLM under structured instructions to verify and refine the graph, ensuring each triplet is visually grounded and temporally consistent. Specifically, following the prompt instruction, the MLLM first removes hallucinated or inconsistent relations, enriches valid events with the coarse- and fine-grained captions. It then applies temporal and logical constraints to ensure consistency: enforcing mutual exclusion of contradictory actions, preserving causal order between prerequisite and consequent events, and maintaining static states across consecutive events unless explicitly terminated in the captions. Formally, the refinement process is defined as:
$\mathcal{G}=f_{\text{MLLM}}(\mathcal{G}_{\text{init}},C^{coarse},C^{fine}),$ (7)
where $\mathcal{G}$ denotes the final verified and enriched Event-based Video Scene Graph. By filtering redundant information and mitigating textual hallucinations in the caption, the event graph provides a more reliable and fine-grained representation of event relations for video reasoning.
Fig. 3 illustrates an example of the Event-based Video Scene Graph (EVSG). An EVSG is composed of event subgraphs derived from multi-grained dense captions. Each event corresponds to a pair of start and end timestamps and includes object nodes and their relations, capturing semantic interactions to model intra-event relations. The event subgraphs are connected through timestamp-based edges that encode their sequential order and temporal dependencies. This hierarchical design enables the EVSG to jointly model intra-event relational reasoning and inter-event temporal reasoning, providing an explicit and interpretable structure that guides the MLLM toward visually grounded and temporally consistent reasoning.
3.4 Event Graph-based Reinforcement Finetuning
Models Param. (B) Type Moment Localization VQA
[…]
GPT-4o [52] - ZS 36.28 45.33 34.00 73.67 28.67
Claude3-Opus [1] - ZS 23.61 30.67 17.67 68.67 13.67
Gemini-1.5-Pro [58] - ZS 28.43 35.67 25.00 68.00 18.33
GPT-4V [51] - ZS 26.74 33.33 22.00 63.33 16.67
Reka-Core [59] - ZS 27.95 36.33 24.00 59.67 17.00
UniVTG [42] - VLP 34.63 53.48 34.53
CG-DETR [47] - VLP 26.53 39.71 22.73
VTimeLLM [25] 7B ZS 20.14 28.84 17.41 36.16 -
TimeChat [56] 7B ZS 11.65 14.42 7.61 40.04 -
LITA [26] 13B ZS 21.49 29.49 16.29 34.44
TOGA  [20] 7B ZS 25.53 29.91 19.79
VTimeLLM [25] 7B SFT 29.92 43.69 26.13 57.58 17.13
TimeChat [56] 7B SFT 26.29 40.13 21.42 49.46 10.92
TimeSearch [53] 7B RL 36.70 48.40 36.40 76.50 29.44
VITAL  [70] 7B RL 40.90 79.10
Qwen2.5-VL [2] 7B ZS 13.60 16.05 9.24 56.60 6.35
GraphThinker(w/o RL) 7B ZS 25.34 33.92 20.25 66.82 15.21
GraphThinker 7B RL 41.46 57.54 40.36 71.30 30.75
[…]
Our GraphThinker integrates EVSGs into the GRPO post-training method, where the graph provides structured event relations that guide and constrain the model’s reasoning process, serving as an intermediate thinking process to reduce hallucinations. The input to the MLLM consists of the video $v$, its corresponding EVSG $\mathcal{G}^{*}$, and a downstream question $q$. The reward design plays a crucial role in shaping the model’s optimization objective within the GRPO framework. To encourage both accurate temporal grounding and interpretable reasoning, we design a composite reward function consisting of three components: a temporal accuracy reward $r_{\text{acc}}(\cdot)$, a format consistency reward $r_{\text{form}}(\cdot)$, and a visual attention reward $r_{\text{attn}}(\cdot)$.
Reward Modeling. (1) Accuracy Reward. The accuracy reward evaluates the quality of the predicted temporal span and the semantic relevance of the generated answer. Specifically, it combines a temporal Intersection over Union (IoU) score with a sentence-level similarity score:
$r_{\text{acc}}(o)=\alpha\,r_{\text{sim}}(o)+(1-\alpha)\,r_{\text{tIoU}}(o),$ (8)
where $r_{\text{tIoU}}$ measures the overlap between the predicted and ground-truth temporal intervals, and $r_{\text{sim}}$ measures semantic similarity between the predicted answer and reference text. This reward encourages the model to reason both temporally and semantically.
(2) Format Reward. To ensure that the model produces interpretable reasoning and structured answers, we introduce a binary template-based reward that verifies whether the prediction adheres to the predefined reasoning format. Specifically, the model must enclose its reasoning within <think>...</think> tags and provide its final grounded response within <answer>...</answer> tags. We apply a regular expression check to determine whether the output adheres to this predefined format. The reward is defined as:
$r_{\text{form}}(o)=\begin{cases}1,&\text{if the $o$ matches the format},\\ 0,&\text{otherwise}.\end{cases}$ (9)
This reward encourages the model to generate well-structured reasoning traces and clearly delimited answers, thereby improving interpretability and stability during training.
(3) Visual Attention Reward. To further encourage the model to balance focus on visual and event-graph cues and mitigate the visual thinking drift problem during reasoning, we introduce a visual attention reward based on the model’s attention distribution. Given the averaged multi-head attention matrix Attn over all token pairs, we define three token groups: textual graph tokens $T_{\text{graph}}$, video tokens $T_{\text{vid}}$, and generated response tokens $T_{\text{res}}$. The attention reward measures the relative proportion of attention that the generated response tokens allocate to visual tokens versus graph tokens:
$r_{\text{attn}}=\frac{\sum\text{Attn}(T_{\text{res}},T_{\text{vid}})}{\sum\text{Attn}(T_{\text{res}},T_{\text{vid}})+\sum\text{Attn}(T_{\text{res}},T_{\text{graph}})}$ (10)
A higher value indicates that the model places more focus on visual evidence rather than abstract graph embeddings, which promotes grounded and perceptually aligned reasoning. This reward is only activated when the reasoning output achieves sufficient semantic and temporal accuracy, specifically when the sentence similarity score exceeds $0.4$ and the temporal IoU score exceeds $0.3$, formulated as:
$r_{\text{attn}}(o)=\begin{cases}r_{\text{attn}}(o),&\text{if }r_{\text{sim}}\geq 0.4\text{ and }r_{\text{tIoU}}\geq 0.3,\\ 0,&\text{otherwise.}\end{cases}$ (11)
(4) Overall Reward. The final reward integrates all components through a weighted combination:
$r(o)=\lambda_{\text{acc}}\,r_{\text{acc}}+\lambda_{\text{form}}\,r_{\text{form}}+\lambda_{\text{attn}}\,r_{\text{attn}},$ (12)
where $\lambda_{\text{acc}},\lambda_{\text{form}},\lambda_{\text{attn}}$ are coefficients controlling the relative contribution of each reward. The overall reward jointly promotes accurate and visually grounded reasoning, allowing the MLLM to effectively leverage the graph for mitigating hallucinations in reasoning.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluated our approach on 2 representative benchmarks that cover event causal reasoning and video hallucination evaluation. (1) Rextime [4] is an event causal relation-based Grounded VQA benchmark. It is designed for complex event reasoning and question answering across time, requiring the model to answer the reasoning question with consistent temporal localization. It contains 9695/921/2143 samples in train/val/test sets. (2) VidHalluc [32] is a benchmark specifically for evaluating video hallucination for MLLM. It consists of 5002 videos with 9295 questions across 3 three critical dimensions for video hallucination evaluation, including Action Hallucination (ACH), Temporal Sequence Hallucination (TSH), and Scene Transition Hallucination (STH).
[…]
Gemini-1.5-Pro [58] $\sim$200B 75.27 79.25 83.83 63.96
GPT-4o [52] $\sim$200B 81.15 90.95 82.00 71.58
Video-ChatGPT [46] 7B 9.50 24.58 30.17 7.70
Video-LLAVA [40] 7B 26.84 64.45 27.17 29.60
ShareGPT4Video [5] 8B 29.96 44.78 49.50 17.83
Chat-UniVi [30] 13B 23.77 54.79 35.50 29.87
LLaVA-NeXT-Video [74] 34B 26.60 77.57 21.33 44.40
PLLaVA [66] 13B 35.30 76.96 16.50 32.44
VideoLLaMA2 [7] 7B 50.04 83.84 26.17 65.12
VILA1.5 [41] 13B 58.46 81.88 63.33 35.03
QwenVL-2.5 [2] 7B 50.46 81.99 67.50 46.77
QwenVL-2.5+EVSG 7B 50.92 83.23 75.33 54.58
GraphThinker(+EVSG+RL)* 7B 66.04 84.57 76.33 57.81
Table 2: Comparisons of multimodal large language models for video reasoning on the three tests in the VidHalluc benchmark, including Action Hallucination (ACH), Temporal Sequence Hallucination (TSH) and Scene Transition Hallucination (STH) test. *denotes finetuning on the Rextime dataset.
Implementation Details. We used QwenVL-2.5-7B [2] as our baseline due to its strong video understanding ability and open-source availability. We also use this baseline for EVSG generation, including generating the multi-grained dense captions and building the EVSG process. The event number limit for dense caption is set to (5, 10, 15). We set $\lambda_{1}{=}0.7$, $\lambda_{2}{=}0.3$, $\lambda_{3}{=}0.6$ and $\alpha{=}0.3$ for the reward function of GRPO post-training. We also utilise the same baseline for event graph generation and input the generated graph during inference. All experiments are conducted using PyTorch on 8 NVIDIA A100 GPUs, with a maximum prompt length of 4096, response length of 2048, 8 rollouts, a batch size of 16, and 1 training epoch.
[…]
4.2 Comparisons of the state-of-the-arts
[…]
VideoHalluc. Tab. 2 shows that our method effectively mitigates the video hallucination problems for the baseline model. When integrating the EVSG to QwenVL-2.5 without any post-training, GraphThinker’s performance across all tasks in the VidHalluc benchmark is notably improved over baseline, especially in the Temporal Sequence Hallucination (TSH) and Scene Transition Hallucination (STH) tasks, improving 7.83% and 7.81%, respectively. These results demonstrate that EVSG’s explicit modeling of event relations in video helps better model understanding in reasoning compared to the implicit event correlation in the baseline. Moreover, GraphThinker also has slight improvement in better action recognition. After conducting GRPO finetuning on Rextime for leveraging EVSG better, GraphThinker achieves the state-of-the-art across the open-source MLLMs. Closed-source models GPT-4o and Gemini-1.5-Pro outperform all open-source models in Tab. 2, partly due to their much larger sizes (GPT-4o: 200B; Gemini-1.5-Pro: 200–600B parameters depending on MoE configuration).
Figure 4: A visual example showing that our method reduces hallucination during reasoning. QwenVL-2.5 still yields hallucination in reasoning during the inference stage. Our method explicitly models event relations to guide the reasoning process for MLLM, producing visually grounded and temporally consistent reasoning.
4.3 Ablation Studies.
We conducted ablation studies with the QwenVL-2.5-3B model on the validation set of the Rextime benchmark to systematically evaluate the contribution of each component in our approach. In Tab. 4.3, we first conduct training-free testing in row 1-3. The results show that our EVSG can benefit the model to understand the event causal relation and achieve more consistent semantic understanding and temporal localization, improving 7.93% on Acc@0.5. Besides, our EVSG demonstrates a better event relation modeling than dense captioning (DC), higher than 3.48% on Acc@0.5. Then, we evaluate post-training testing. Applying GRPO alone yields moderate improvements, but combining it with our EVSG leads to significant gains across all metrics. Notably, the integration of the $r_{attn}$ in GRPO further achieves better performance, showing that our proposed $r_{attn}$ better leverages the EVSG and high-value visual cues in videos for reasoning.
Methods mIoU Acc Acc@0.5
[…]
QwenVL-2.5-7B 13.77 55.65 6.20
QwenVL-2.5-7B + DC 24.25 62.87 15.42
QwenVL-2.5-7B + EVSG 25.36 63.30 16.07
[…]
QwenVL-2.5-7B 41.04 57.55 27.03
QwenVL-2.5-7B + DC 41.20 69.60 29.10
QwenVL-2.5-7B + DC + $r_{attn}$ 41.42 71.01 32.72
QwenVL-2.5-7B + EVSG 42.57 72.20 32.03
QwenVL-2.5-7B + EVSG + $r_{attn}$ 43.56 73.72 33.49
Table 3: Ablation study of QwenVL-2.5-7B with various training strategies on the validation set of the Rextime benchmark. “DC” denotes dense captioning.
Event Numbers mIoU Acc Acc@0.5
(3,5,7) 26.02 63.84 16.94
(5,10,15) 27.98 64.17 18.02
(10,15,20) 25.35 64.06 16.83
Table 4: Parameter analysis of QwenVL-2.5-7B+EVSG under the training-free setting on the Rextime validation set. The event number limit controls the maximum number of events generated in dense captioning for event graph generation.
Tab. 4 analyzes the impact of using different event number limits in multi-grained caption generation for event graph construction. In our experiment, most of our videos are around 3 minutes long. The results indicate that too few events lead to captions that fail to capture sufficient video details, while too many events make the captions overly verbose, introducing redundancy and hallucinations. A moderate number of events (5, 10, 15) makes a better balance between detail and redundancy, resulting in more coherent and informative event graphs. We provide more parameter analysis in the supplementary materials.
4.4 Qualitative Analysis
Fig. 4 provides a visual example. It shows that our method improves object and event relation modeling, enhances event localization, and reduces hallucinations during multi-step reasoning. In this example, previous methods incorrectly interpret the scene as the man performing a dynamic warm-up, leading to the wrong answer (B). In contrast, our EVSG-enhanced model correctly identifies the detailed temporal sequence: the man first performs sit-ups on the bench, then gets off, folds a white towel on the ground, and finally begins rolling crunches with the ab wheel. By explicitly grounding these actions in the temporal context, our method produces the correct answer (C) with precise temporal localization ([0-49.1s]). This qualitative case illustrates that our method allows the model to better capture causal and sequential dependencies between events, yielding more consistent reasoning and alleviating hallucinations caused by implicit event relations.
5 Conclusion
This work introduces GraphThinker, a MLLM post-training method integrated with explicit event-level graphs for video reasoning, effectively reducing hallucinations during multi-step reasoning. To explicitly model intra-event and inter-event relations without manual annotations, we propose an Event-based Video Scene Graph (EVSG) generated by the MLLM itself. For better leveraging the event-level graph for video reasoning, we integrate it into the GRPO post-training framework and design a visual attention reward to effectively utilize EVSG and high-value visual cues. Our method shows strong effectiveness and achieves the state-of-the-art on Rextime and VidHalluc benchmarks.

### v2

@@
 This often leads to hallucinations in video reasoning.
-For instance, when determining the temporal order of events, QwenVL-2.5 tends to yield temporal hallucinations due to a lack of explicit event-level evidence validation and reasoning.
+For instance, when determining the temporal order of events, Qwen2.5-VL tends to yield temporal hallucinations due to a lack of explicit event-level evidence validation and reasoning.
 In contrast, our GraphThinker explicitly models both intra- and inter-event relations in accordance with selective visual attentions in a video to constrain reasoning under these structured event relations anchored by fine-grained visual evidence, reducing hallucinations and improving temporal consistency in reasoning.
@@
 VILA1.5 [41] 13B 58.46 81.88 63.33 35.03
-QwenVL-2.5 [2] 7B 50.46 81.99 67.50 46.77
-QwenVL-2.5+EVSG 7B 50.92 83.23 75.33 54.58
+Qwen2.5-VL [2] 7B 50.46 81.99 67.50 46.77
+Qwen2.5-VL+EVSG 7B 50.92 83.23 75.33 54.58
 GraphThinker(+EVSG+RL)* 7B 66.04 84.57 76.33 57.81
@@
 Implementation Details.
-We used QwenVL-2.5-7B [2] as our baseline due to its strong video understanding ability and open-source availability.
+We used Qwen2.5-VL-7B [2] as our baseline due to its strong video understanding ability and open-source availability.
 We also use this baseline for EVSG generation, including generating the multi-grained dense captions and building the EVSG process.
@@
 2 shows that our method effectively mitigates the video hallucination problems for the baseline model.
-When integrating the EVSG to QwenVL-2.5 without any post-training, GraphThinker’s performance across all tasks in the VidHalluc benchmark is notably improved over baseline, especially in the Temporal Sequence Hallucination (TSH) and Scene Transition Hallucination (STH) tasks, improving 7.83% and 7.81%, respectively.
+When integrating the EVSG to Qwen2.5-VL without any post-training, GraphThinker’s performance across all tasks in the VidHalluc benchmark is notably improved over baseline, especially in the Temporal Sequence Hallucination (TSH) and Scene Transition Hallucination (STH) tasks, improving 7.83% and 7.81%, respectively.
 These results demonstrate that EVSG’s explicit modeling of event relations in video helps better model understanding in reasoning compared to the implicit event correlation in the baseline.
@@
 Figure 4: A visual example showing that our method reduces hallucination during reasoning.
-QwenVL-2.5 still yields hallucination in reasoning during the inference stage.
+Qwen2.5-VL still yields hallucination in reasoning during the inference stage.
 Our method explicitly models event relations to guide the reasoning process for MLLM, producing visually grounded and temporally consistent reasoning.
 4.3 Ablation Studies.
-We conducted ablation studies with the QwenVL-2.5-3B model on the validation set of the Rextime benchmark to systematically evaluate the contribution of each component in our approach.
+We conducted ablation studies with the Qwen2.5-VL-3B model on the validation set of the Rextime benchmark to systematically evaluate the contribution of each component in our approach.
 In Tab.
@@
 […]
-QwenVL-2.5-7B 13.77 55.65 6.20
-QwenVL-2.5-7B + DC 24.25 62.87 15.42
-QwenVL-2.5-7B + EVSG 25.36 63.30 16.07
+Qwen2.5-VL-7B 13.77 55.65 6.20
+Qwen2.5-VL-7B + DC 24.25 62.87 15.42
+Qwen2.5-VL-7B + EVSG 25.36 63.30 16.07
 […]
-QwenVL-2.5-7B 41.04 57.55 27.03
-QwenVL-2.5-7B + DC 41.20 69.60 29.10
-QwenVL-2.5-7B + DC + $r_{attn}$ 41.42 71.01 32.72
-QwenVL-2.5-7B + EVSG 42.57 72.20 32.03
-QwenVL-2.5-7B + EVSG + $r_{attn}$ 43.56 73.72 33.49
-Table 3: Ablation study of QwenVL-2.5-7B with various training strategies on the validation set of the Rextime benchmark. “DC” denotes dense captioning.
+Qwen2.5-VL-7B 41.04 57.55 27.03
+Qwen2.5-VL-7B + DC 41.20 69.60 29.10
+Qwen2.5-VL-7B + DC + $r_{attn}$ 41.42 71.01 32.72
+Qwen2.5-VL-7B + EVSG 42.57 72.20 32.03
+Qwen2.5-VL-7B + EVSG + $r_{attn}$ 43.56 73.72 33.49
+Table 3: Ablation study of Qwen2.5-VL-7B with various training strategies on the validation set of the Rextime benchmark. “DC” denotes dense captioning.
 Event Numbers mIoU Acc Acc@0.5
@@
 (10,15,20) 25.35 64.06 16.83
-Table 4: Parameter analysis of QwenVL-2.5-7B+EVSG under the training-free setting on the Rextime validation set.
+Table 4: Parameter analysis of Qwen2.5-VL-7B+EVSG under the training-free setting on the Rextime validation set.
 The event number limit controls the maximum number of events generated in dense captioning for event graph generation.